the height of an above ground pool is 3 feet. the pool needs to be drained. as the water drains, the height of the water changes at a rate of -(1)/(2) inch per minute. write and sold and equation to find how many minutes it will take to drain the pool
1
Expert's answer
2019-04-03T10:11:06-0400

3 ft=3*12=36 in.

36-(1/2)x=0

x=36*2=72 min.